Paper wasps are a haplodiploid, eusocial species. In some nests, there is a secondary foundress (queen helper). Only The primary foundress produces offspring. The secondary foundress is a sister of the primary foundress in 50% of the nests and completely unrelated in the other 50% of nests. In one study of paper wasps, nests without helpers produced 40 reproductive offspring while nests with secondary foundress produced 80 reproductive offspring. Answer the following questions based off this information.

A.) What is the coefficient of relatedness between the eusocial sister wasps?

B.) Keeping in mind that the benefit of altruism (B) is measured by the increase in reproductive success brought about by the altruistic behavior, calculate B for these paper wasps.

C.) The cost of altruism (C) is measured as the difference between the number of offspring produced by an unhelped primary foundress and the number produced directly by the secondary foundress. Calculate C.

D.) Calculate the average degree of relatedness (coefficent of relatedness) between the secondary and the primary foundresses. (HINT 1/2 of the secondary foundresses are sisters, the other 1/2 are completely unrelated)

E.) Use Hamilton's rule (Br-C>o) to discuss whether this altruistic behavior could have evolved through kin selection alone.  
Answer #1

A) Eusocial sister wasps are diploid since both are female. Females arise from fertilized eggs.(diploid) Mother (haploid) Father 1/2 Sister Sister

Relatedness between eusocial sister wasps  (+1) | +12

D) Relatedness between unrelated female wasps = 0

Therefore, relatedness between primary foundress and secondary foundress  | |

B) Benefit of altruism  = 80-40 = 40

C) Cost of altruism  = 40 -0= 40

E)

Bxr-C = 40 x - 40 < 0

According to Hamilton's rule, an individual organism will show altruistic behavior to favour indirect fitness over direct fitness if it increases its inclusive fitness. It happens when the actor and reciepent of altruism are related by blood. This type of natural selection favoring indirect fitness is called kin selection.

Inclusive Fitness = Direct Fitness + Indirect Fitness

Hamilton's rule:  Вхт - с 0

Since the above situation in paper wasps do not follow Hamilton's rule, the actor do not have an increase in its inclusive fitness by showing altruistic behavior. Therefore, this altruistic behavior could not have evolved through kin selection alone.